OpenStudy (anonymous):

Find all the solutions to the equation (check all that apply). 20x^2-26x+8=0 a. x=1/3 b. x=1/5 c. x=3/5 d. x=1/2 e. x=2/3 f. x=4/5

OpenStudy (anonymous):

substitute the values of x in the equation

OpenStudy (anonymous):

= 2(10x^2 - 13x + 4) = 0 (5x - 4 )( 2x - 1) = 0 are u ok from here?

OpenStudy (anonymous):

Yeah I got it thanks jimmy!

OpenStudy (anonymous):

d & f